OpenStudy (anonymous):

what two numbers multiplied together get -48 but when added together = 10

OpenStudy (anonymous):

Solve with a system: \[a\cdot b = -48\]\[\implies a = -\frac{48}{b}\]\[a+b = 10\]\[\implies -\frac{48}{b} + b = 10\]\[\implies -48 + b^2 = 10b\]\[\implies b^2 - 10b - 48 = 0\]\[\implies b = 5 \pm \sqrt{73}\] Whichever one of these two you want b to be, a will be the other one.

jhonyy9 (jhonyy9):

not exists like numbers